# Modeling scale-up of particle coating by atomic layer deposition

### v1

@@ I Introduction
One of the perceived drawbacks of ALD is that it is a slow process: in addition to its pulsed, sequential nature – with each ALD cycle involving alternate exposures to two different precursors separated by purge times, in self-limited processes the reactivity of the surface goes down as the number of available reaction sites decreases. This typically reduces throughput compared to non self-limited processes like chemical vapor deposition (CVD).
Intuitively, we understand that as we scale an ALD process to larger batch sizes and increasingly higher surface areas, the process will eventually transition from a reaction limited regime to a transport limited regime, thereby eliminating some of the inefficiencies related to the self-limited surface kinetics. Currently, though, we do not have a good understanding of when and how the transition takes place, and how the underlying surface kinetics and the manufacturing approach affects this transition. This complicates the transfer of ALD processes from the lab to manufacturing.
In this work, we explore the fundamentals of the scale-up of particle coating by ALD. In particular, we focus on three common manufacturing approaches for particle coating: rotating drumMcCormick et al. (2007); Coile et al. (2020), fluidized bedHakim et al. (2006), and continuous reactorsvan Ommen et al. (2015); Hartig et al. (2021), generalizing traditional chemical engineering models to incorporate the spatial and time dependencies intrinsic to self-limited processesYanguas-Gil and Elam (2012a). In a prior work, we showed how this approach provided excellent qualitative and quantitative agreement for the specific case of a fluidized bed reactorLu et al. (2022). There is therefore an opportunity to apply the same methodology to other scaling approaches and to expand the range of surface kinetics models considered.
@@ II.1.1 Ideal self-limited model
A key feature of self-limited processes is the presence of a finite number of reactive sites on the surface. Using the open site formalismKee et al. (2018), the surface reactivity can be represented by a sticking probability, $\beta$, that depends on the fraction of surface sites available for precursor molecules to react with at a given time. If we define the fractional coverage, $\Theta$, as the fraction of surface sites that have already reacted, the simplest model introduces a first-order dependence on the fraction of available sites, $1-\Theta$:
@@ II.1.2 Extension to heterogeneous surfaces
The model above represents the ideal case where the surface can be partitioned into identical reaction sites. However, in many cases of interest surfaces are heterogeneous, which means that we need to extend Eq. 1 to consider the presence of multiple sites:
$\beta=\sum_{s}f_{s}\beta_{s}\left(1-\Theta^{s}\right)$ (4)
With $f_{s}$ representing the fraction of each site, $\beta_{s}$ is the reaction probability of the precursor molecule with that site, and $\sum_{s}f_{s}\leq 1$ is the total fraction of reactive surface sites. There are two cases of interest for this work: the first is when only a fraction of the surface is reactive towards the ALD precursor. In this case, if $f$ represents the fraction of the sites that are reactive, we have that:
@@
resulting in a lower sticking probability, with the evolution of $\Theta$ still given by Eq. 2. This case is representative of surface functionalization or passivation approaches where nucleation is potentially sparse. Including $f$ in our model allows us to independently tune the size of a surface site, given by its area $s_{0}$, and the number of active sites on the surface. The second case comprises surfaces with two types of reactive sites, one with high and another one with low sticking probabilityBielinski et al. (2022). This scenario allows us to consider “soft-saturating” ALD processes, where after an initially fast mass uptake, the growth per cycle as a function of dose time asymptotically reaches a limiting valueYanguas-Gil, Libera, and Elam (2021). In this scenario, the sticking probability is given by:
@@
In this case, both $\Theta_{a}$ and $\Theta_{b}$ evolve as a function of time following Eq. 2.
@@ II.2 Reactor models
In this work we introduce models for two batch and two continuous ALD particle coating strategies (Table 1). For the batch cases, which correspond to rotating drum and fluidized bed reactors, we assume homogeneous particle mixing, whereas for the continuous processes we assume that mixing takes place preferentially in the directions perpendicular to particle transportHartig et al. (2021, 2023). Examples of continuous processing include continuous vibrating reactors (CVR). A second key assumption involves how precursor delivery takes place in these reactors: the well-mixed approximation is a valid assumption when the geometry of the reactor favors mixing and mixing takes place faster than the characteristic dose time [Fig. 1(A)]. Alternatively, the reactor can be designed to force precursor to flow through the agitated particles. For fluidized bed ALD reactors, in a prior work we showed that a plug flow model of precursor transport reproduced well reactor configurations where precursor is introduced upstream of the reactor and is delivered by the same flow used for particle fluidizationLu et al. (2022) [Fig. 1(B)]. Finally, for the continuous fluidized bed we consider both the well-mixed and cross-flow cases, which would correspond to short and long reactor geometries [Figs. 1(C) and 1(D), respectively]. The mapping between these models and strategies for particle coatings is not one-to-one: for instance, fluidized bed reactors can be designed to operate in batch or continuous mode; rotating drum reactors can have geometries that favor mixing or that transport the precursor along the axis of revolution. Some of the possible combinations are enumerated in Table 1.
@@ D Continuous In plane mixing Plug flow CVR, fluidized bed, spatial ALD
These four models share a series of common parameters, including the reactor volume, $V$, total surface area to be coated, $S$, and the precursor residence time, $t_{\mathrm{res}}$. Also, $n_{0}$ and $p_{0}$ are the precursor molecular density and pressure at the inlet, respectively. Models that explicitly consider either particle or gas transport in the reactor depend on the reactor length, $L$, its cross-sectional area, $S_{0}$, the flow velocity, $u$, and the particle velocity, $v$.
@@ II.2.1 Batch process with well-mixed reactor approximation
This model assumes ideal mixing so that all particles are homogeneously coated [Figure 1(A)]. Furthermore, the precursor is also efficiently mixed in the reactor, so that the well-mixed approximation applies. An example would be a rotating drum reactor under near-static precursor flow conditions.
@@
Whereas for a soft-saturating system the model equations are:
@@
This model is representative of particle coating approaches such as rotating drum reactors.
@@ II.2.2 Batch process with plug flow precursor transport
We next consider a plug flow model for the precursor transport across a homogenously mixed particle column of cross section area $S_{0}$ and length $L$ [Figure 1(B)]. Examples for this type of reactor include a fluidized bed or a rotating drum under cross-flow conditions. In this case, the precursor balance is given by:
@@
where $u$ is the flow velocity. This leads to the following equations for the single reaction pathway case:
@@
where $\bar{n}$ is the average precursor density along the column.
@@ II.2.3 Continuous process and well-mixed precursor approximation
We next consider particles moving with a velocity $v$ inside a reactor of length $L$ [Figure 1(C)]. The particles are only partially mixed, with perfect mixing only in the plane perpendicular to the direction of movement of the particles. Consequently, the surface coverage depends only on the particle position along the upstream-downstream axis: $\Theta=\Theta(z)$ and on the total exposure received until reaching $z$. Furthermore, the well-mixed approximation holds for the ALD precursor transport. An example for this type of system would be a vibro-fluidized bed reactor under near static flow conditions. For a single heterogeneous process, the model equations are:
@@ II.2.4 Continuous process with plug flow precursor transport
The final model considers the case of a continuous process in which the precursor is injected in an upstream position and flows downstream over a fluidized bed of particles moving in the same direction [Figure 1(D)]. In this case, both the fractional surface coverage and the precursor concentration depend on the axial position. For the case of a single heterogeneous process:
@@ II.3 Model normalization
In order to establish a comparison between these models we will use two common non-dimensional numbers: the first is the Damköhler number, which can be defined as the ratio of the characteristic reaction rate and the transport rate:
@@
The resulting non-dimensional equations for the four models in both the single heterogeneous process case and the soft-saturation case are provided in the Supporting Information.
@@ III.2 Scale up of batch processes
The value of $\mathrm{Da}$ increases with increasing amount of surface area inside the reactor. Consequently, it is a useful parameter that can be applied to track the behavior of a process upon scale up. An ideal process would achieve saturation in the minimum possible time and utilize 100% of the supplied precursor. As will be shown below, all of the reactor models approach ideal behavior in the limit of high $\mathrm{Da}$, but the degree of ideality varies among the models and this highlights differences in their scalability.
Figure 2 shows saturation plots with the evolution of the fractional surface coverage as a function of dose time for batch processes with the well-mixed [Fig 2(A)] and plug flow [Fig 2(B)] precursor transport approximations. Results are shown for increasing values of the Damköhler number.
@@
The plots are represented using the normalized dose time $\tau$. A value of $\tau=1$ represents the time at which the total number of moles of precursor inserted in the reactor equals the total number of surface reactive sites, so a fractional surface coverage approaching one for $\tau=1$ indicates that the process is reaching the transport-limited regime and approaching ideal behiavior. For both the well-mixed and plug flow models, the slope of fractional coverage vs dose time increases with increasing values of $\mathrm{Da}$. The time to saturation also shortens, reaching values close to $\tau=1$, thus confirming that ALD processes do become more efficient as they are scaled. However, that transition takes place at lower Damköhler numbers for the plug flow model, with the transition taking place for $\mathrm{Da}>10$, whereas for the well-mixed reactor the transition requires $\mathrm{Da}>30$.
@@
A crucial metric for manufacturing is precursor efficiency, which we define as the percentage of precursor molecules that are used in the deposition process. In Figure 4 we show the precursor efficiency for both the well-mixed and plug flow models as a function of the final fractional surface coverage for increasing values of the Damköhler number. At low Damköhler numbers, the precursor utilization is low due to the intrinsic slow down of the surface kinetics in self-limited processes. However, as the Damköhler number increases, the process becomes increasingly more efficient until reaching almost 100% precursor utilization. As shown above for the fractional surface coverage, the plug flow model yields a higher precursor efficiency than the well-mixed model for identical fractional surface coverage and Damköhler number.
@@
Likewise, in Figure 5 we show the fraction of unreacted precursor that leaves the reactor as a function of normalized dose time for both batch processes. This depiction of precursor efficiency is useful because it models a downstream measurement of the precursor concentration using a mass spectrometer. The high precursor efficiencies shown in Figure 4 consistently lead to almost no precursor leaving the reactor prior to reaching saturation. However, as the system transitions to a transport limited regime, the “punch-through” signal of unreacted precursor becomes increasingly more sharp at the point where saturation is reached. This agrees well with experimental observations showing how in-situ mass spectrometry provides a characteristic punch-through signal for particle coated by a TMA/H2O ALD process in a fluidized bed reactorWeimer (2019); Lu et al. (2022).

### v2

@@ I Introduction
One of the perceived drawbacks of ALD is that it is a slow process: first, ALD is a pulsed, sequential process, with each ALD cycle involving alternate exposures to two different precursors separated by purge times. Second, surfaces in self-limited processes become less reactive with time as the number of available surface sites decreases. These two factors tend to reduce ALD’s throughput compared to non self-limited processes like chemical vapor deposition (CVD).
Intuitively, we understand that, as we scale an ALD process to larger batch sizes and increasingly higher surface areas, the process will eventually transition from a reaction limited to a transport limited regime, thereby eliminating some of the inefficiencies related to the self-limited surface kinetics. Currently, though, we do not have a good understanding of when and how this transition takes place or how it is affected by either the process surface kinetics or the manufacturing approach. This complicates the transfer of ALD processes from the lab to large scale manufacturing.
In this work, we explore the fundamentals of the scale-up of particle coating by ALD. In particular, we focus on three common manufacturing approaches for particle coating: rotating drumMcCormick et al. (2007); Duan et al. (2015); Coile et al. (2020), fluidized bedHakim et al. (2006), and continuous reactorsvan Ommen et al. (2015); Hartig et al. (2021). As the interest in ALD manufacturing has increased, a number of works have tackled particle coating in ALD from a simulation standpointGrillo, Kreutzer, and van Ommen (2015); Duan et al. (2017); Jin, Kleijn, and van Ommen (2017); Hartig et al. (2023). These works have usually focused on a single specific particle coating technique. Our approach is to generalize traditional chemical engineering models to incorporate the spatial and time dependencies intrinsic to self-limited processesYanguas-Gil and Elam (2012a). This allows us to compare the scaling behavior of different ALD particle coating strategies. In a prior work, we showed how in the specific case of a fluidized reactor such simple models provide excellent qualitative and quantitative agreement with experimental dataLu et al. (2022). There is therefore an opportunity to apply the same methodology to other scaling approaches and to expand the range of surface kinetics models considered.
@@ II Model
In this section we introduce the models for the ALD of particles on rotating drum, fluidized bed, and continuous reactors. In Section II.1 we describe the surface kinetic models, while in Section II.2 we introduce the different reactor models. A list of the symbols used is presented in Table 1.
@@ II.1.1 Ideal self-limited model
A key feature of self-limited processes is the presence of a finite number of reactive sites on the surface. Using the open site formalismKee et al. (2018), the surface reactivity can be represented by a sticking probability, $\beta$, which depends on the fraction of surface sites available for precursor molecules to react with at a given time. If we define the fractional coverage, $\Theta$, as the fraction of surface sites that have already reacted, the simplest model introduces a first-order dependence on the fraction of available sites, $1-\Theta$:
@@
Table 1: List of symbols used in this work
Symbol Interpretation
$\beta$ total precursor sticking probability
$\beta_{i}$ reaction probability with an available site $i$
$f_{i}$ fraction of surface site $i$
$\Theta_{i}$ fractional surface coverage of surface site $i$
$s_{0}$ average area of a surface site
$J$ surface flux of precursor molecules
$v_{th}$ precursor mean thermal velocity
$p$ precursor partial pressure
$n$ precursor molecular density
$n_{0}$ precursor molecular density at inlet
$T$ process temperature
$V$ reactor volume
$S$ total particle surface area
$t_{\mathrm{res}}$ precursor residence time
$u$ flow velocity in plug flow model
$S_{0}$ reactor cross-section area in plug flow model
$L$ reactor length in plug flow model
$v$ velocity of moving bed of particles
$\phi$ carrier gas volumetric flow
$t_{0}$ time at which the number of precursor molecules inserted
in the reactor is equal to the number of surface sites
$t_{s}$ particle residence time in a continuous process
$\mathrm{Da}$ Damköhler number
$\gamma$ excess number, precursor molecule to surface site ratio
$\tau$ normalized dose time for batch models
$\tau_{s}$ normalized particle residence time for continuous process
$\xi$ normalized reactor length for plug flow models
$x$ normalized precursor concentration
@@ II.1.2 Extension to heterogeneous surfaces
The model above represents the ideal case where the surface can be partitioned into identical reaction sites. However, in many cases of interest surfaces are heterogeneous, and we need to extend Eq. 1 to consider the presence of multiple sites:
$\beta=\sum_{s}f_{s}\beta_{s}\left(1-\Theta_{s}\right)$ (4)
Here $f_{s}$ represents the fraction of each site, $\beta_{s}$ is the reaction probability of the precursor molecule with that site, and $\sum_{s}f_{s}\leq 1$ is the total fraction of reactive surface sites. In this work we consider the following two cases: in the first scenario only a fraction of the surface is reactive towards the ALD precursor. In this case, if $f$ represents the fraction of the sites that are reactive, we have that:
@@
resulting in a lower sticking probability, with the evolution of $\Theta$ still given by Eq. 2. This scenario is representative of surface functionalization or passivation approaches where nucleation is potentially sparse. Including $f$ in our model allows us to independently tune the size of a surface site, given by its area $s_{0}$, and the number of active sites on the surface.
The second scenario comprises surfaces with two types of reactive sites, one with high and another with low sticking probabilityBielinski et al. (2022). This scenario allows us to consider “soft-saturating” ALD processes where, after an initially fast mass uptake, the growth per cycle as a function of dose time asymptotically reaches a limiting valueYanguas-Gil, Libera, and Elam (2021). In this scenario, the sticking probability is given by:
@@
In this case, both $\Theta_{a}$ and $\Theta_{b}$ evolve as a function of time following Eq. 2. If $f_{a}+f_{b}=1$, we can make $f_{b}=f$ and Eq. 6 can be expressed as:
$\beta=(1-f)\beta_{a}(1-\Theta_{a})+f\beta_{b}(1-\Theta_{b})$ (8)
@@ II.2 Reactor models
There are two key characteristics that we can use to classify different particle coating strategies. We will use them to build four different models for ALD particle coating that cover all possible combinations (Fig. 1).
A first distinction is whether particles are coated in batches or are continuously fed into the reactor. Examples of batch processes include rotating drum and fluidized bed reactors. For the batch models, we assume homogeneous particle mixing. This corresponds to experimental configurations in which the characteristic dose time is much larger than the characteristic particle mixing time. In this regime, all particles are homogeneously coated as they sample conditions all over the reactor.
In continuous processes, particles are fed into the reactor continuously, being exposed to the ALD precursor as they travel inside the reactor [Figs. 1(C) and (D)]. Examples of continuous processes include continuous vibrating reactors (CVR) and continuous fluidized reactors. For continuous processes, we assume that particles are fluidized or agitated in a way that achieves perfect mixing in directions perpendicular to particle transportHartig et al. (2021, 2023). As a consequence, the surface coverage of particles increases as particles move further along the reactor. The velocity of the moving particles $v$ and the length of the reactor $L$ are key variables in continuous processes, as they determine the particle residence time inside the reactor. Due to the perfect mixing, particles exiting the reactor all have identical surface coverage. In both the batch and continuous cases, a narrow distribution of film thickness is a necessary condition for these approximations to hold.
A second distinction involves precursor transport inside the reactor: when the geometry and configuration of the reactor favors precursor mixing, the precursor mixing time is much faster than the characteristic dose time. Under these conditions, we can model precursor transport using the well-mixed approximation [Fig. 1(A)], and the reactor volume $V$ and characteristic residence time $t_{\mathrm{res}}$ are the key variables defining precursor transport. Rotating drum reactors and configurations that force precursor mixing are examples of processes where the well-mixed approximation can hold.
Alternatively, the reactor can be designed to force precursor to flow through it. This is the case of of some fluidized bed reactors with tall and narrow fluidized beds [Fig. 1(B)]. Prior works involving ALD on both cross-flow reactors and particle coating in fluidized bed reactors have shown that plug flow models of precursor transport reproduced well the experimental resultsLu et al. (2022); Yanguas-Gil and Elam (2014). In reactors where precursor transport can be approximated by a plug flow model, the flow velocity $u$, the cross-sectional area of the reactor $S_{0}$, and the reactor length $L$ are key variables.
The well-mixed and plug flow approximations can be viewed as two limiting cases for precursor transport that can be applied to both batch and continuous ALD processes. Consequently, in this work we explore all four possible combinations, as shown in Fig. 1 and Table 2. We should emphasize that the mapping between these four models and strategies for particle coatings is not one-to-one: for instance, depending on the reactor geometry, continuous vibrating reactors can be modeled using the well-mixed and the plug flow approximations [Figs. 1(C) and 1(D), respectively]. Likewise, fluidized bed reactors can be designed to operate in batch or continuous mode, and rotating drum reactors can have geometries that favor mixing or that transport the precursor along the axis of revolution. Some of the possible combinations are enumerated in Table 2.
@@ D Continuous In plane mixing Plug flow CVR, fluidized bed, spatial ALD
In all these cases, the particle loading is determined by the total surface area $S$ of the particles inside the reactor. When particle loadings are high, particles can occupy a non-negligible fraction $\varepsilon$ of the particle volume $V$ and the reactor cross-sectional area $S_{0}$. These factors can be incorporated into the model constants. Finally, the number of precursor moles inserted in the reactor is proportional to the precursor molecular density $n_{0}$ or, equivalently to the precursor partial pressure $p_{0}$, at the inlet. These will be used as initial conditions in all the models considered in this work.
@@ II.2.1 Batch process with well-mixed reactor approximation
This model assumes ideal mixing so that all particles are homogeneously coated [Figure 1(A)]. Under the assumption that all particles are homogeneously mixed, we only need to consider a single value for the fractional surface coverage. Furthermore, the precursor is also efficiently mixed in the reactor, and the well-mixed approximation applies. A prototypic example would be a rotating drum reactor under near-static precursor flow conditions. As mentioned in Section II.2, the precursor transport is defined by a residence time $t_{\mathrm{res}}$.
@@
Whereas for a soft-saturating system (Eq. 8) the model equations are:
@@ II.2.2 Batch process with plug flow precursor transport
We next consider a plug flow model for the precursor transport across a homogeneously mixed particle column of cross section area $S_{0}$ and length $L$ [Figure 1(B)]. Examples for this type of reactor include a fluidized bed or a rotating drum reactor under cross-flow conditions. In this case, the precursor transport is modeled using a plug flow model:
@@
where $u$ is the flow velocity and $z$ is the distance from the precursor inlet.
Under the assumption that the characteristic time for particle mixing is much smaller than the dose times required to achieve saturation, particles are equally likely to be at any position within the reactor during a precursor dose. Consequently, they will be exposed to the average precursor concentration in the reactor at any given time, and the surface kinetics will depend on the average precursor density $\bar{n}$:
[…]
This leads to the following equations for the single reaction pathway case:
@@ II.2.3 Continuous process and well-mixed precursor approximation
We next consider the case of a continuous process. As described in Section II.2, we assume that particles move with a velocity $v$ inside a reactor of length $L$ [Figure 1(C)]. The particles are only partially mixed, with perfect mixing assumed only in the plane perpendicular to the direction of movement. Consequently, the surface coverage will depend on the particle position along the upstream-downstream axis $z$, $\Theta=\Theta(z)$, and on the total exposure received until reaching $z$. Furthermore, in this model we assume that the precursor is well-mixed inside the reactor. An example for this type of system would be a vibro-fluidized bed reactor under near static flow conditions.
Under the well-mixed approximation, the precursor density is given by a balance between the moles of precursor coming into the reactor at any given time and the precursor losses due to transport out of the reactor and the heterogeneous processes. The total precursor loss to surface reactions can be calculated by integrating the sticking probability $\beta$ along the upstream-downstream axis.
For a single heterogeneous process, this leads to the following equations:
@@
This model is very similar to a model we previously developed for spatial ALDYanguas-Gil and Elam (2014), except that in this case the surface area is dominated by the total particle surface area $S$.
@@ II.2.4 Continuous process with plug flow precursor transport
The final case considered in this work comprises a continuous process in which the precursor is injected in an upstream position and flows downstream over a fluidized bed of particles moving in the same direction [Figure 1(D)]. In this case, both the fractional surface coverage and the precursor concentration depend on the axial position $z$. For the case of a single heterogeneous process:
@@
Here $v$ is the velocity of the particles, as defined for the continuous well-mixed model, $S_{0}$ is the cross-sectional area of the reactor and $u$ is the flow velocity.
@@ II.3 Model normalization
In order to establish a comparison between the models introduced in Section II.2, we will use two common non-dimensional numbers: the first is the Damköhler number, which can be defined as the ratio of the characteristic reaction rate and the transport rate:
@@
The resulting non-dimensional equations for the four models in both the single heterogeneous process case and the soft-saturation case are provided in the Appendix.
@@ III.2 Scale up of batch processes
The value of $\mathrm{Da}$ increases with increasing amount of surface area inside the reactor. Consequently, we can use this parameter to track the behavior of a process upon scale up. An ideal process would achieve saturation in the minimum possible time and utilize 100% of the supplied precursor. As will be shown below, all of the reactor models approach ideal behavior in the limit of high $\mathrm{Da}$, but the degree of ideality varies among the models. This highlights the differences in their scalability.
@@
Figure 2 shows saturation plots depicting the evolution of the fractional surface coverage as a function of dose time for batch processes with the well-mixed [Fig 2(A)] and plug flow [Fig 2(B)] precursor transport approximations. Results are shown for increasing values of the Damköhler number. The plots are represented using the normalized dose time $\tau$. A value of $\tau=1$ represents the time at which the total number of moles of precursor inserted in the reactor equals the total number of surface reactive sites, so a fractional surface coverage approaching one for $\tau=1$ indicates that the process is reaching the transport-limited regime and approaching ideal behavior. For both the well-mixed and plug flow models, the slope of fractional coverage vs dose time increases with increasing values of $\mathrm{Da}$. The time to saturation also shortens, reaching values close to $\tau=1$, thus confirming that ALD processes do become more efficient as they are scaled. However, that transition takes place at lower Damköhler numbers for the plug flow model, with the transition taking place for $\mathrm{Da}>10$, whereas for the well-mixed reactor the transition requires $\mathrm{Da}>30$.
@@
A crucial metric for manufacturing is precursor efficiency, which we define as the percentage of precursor molecules that are used in the deposition process. In Figure 4 we show the precursor efficiency for both the well-mixed [Fig. 4(A)] and plug flow [Fig. 4(B)] models as a function of the final fractional surface coverage for increasing values of the Damköhler number. At low Damköhler numbers, the precursor utilization is low due to the intrinsic slow down of the surface kinetics in self-limited processes. However, as the Damköhler number increases, the process becomes increasingly more efficient until reaching almost 100% precursor utilization. As shown above for the fractional surface coverage, the plug flow model yields a higher precursor efficiency than the well-mixed model for identical fractional surface coverage and Damköhler number.
@@
Likewise, in Figure 5 we show the fraction of unreacted precursor that leaves the reactor as a function of normalized dose time for both batch processes. This depiction of precursor efficiency is useful because it models a downstream measurement of the precursor concentration using a mass spectrometer. The high precursor efficiencies shown in Figure 4 consistently lead to almost no precursor leaving the reactor prior to reaching saturation. However, as the system transitions to a transport limited regime, the “punch-through” signal of unreacted precursor becomes increasingly more sharp at the point where saturation is reached.
The model results agree well with experimental observations on fluidized bed reactors: first, in-situ mass spectrometry provides a characteristic punch-through signal similar to that shown in Fig. 5Weimer (2019); Lu et al. (2022). In a prior work focused on Al2O3 ALD from TMA/H2O ALD, we also observed a similarity between the precursor pressure traces observed at the inlet and the mass spectrometry signals from reaction byproducts in the downstream positionLu et al. (2022). This is consistent with the 100% precursor efficiency expected at high Damköhler numbers (Fig. 4). Finally, in that same work we observed that the saturation times, defined as the time at which punch-through was experimentally observed, also agreed well with the predictions from Eq. 41. This indicates that, despite their simplicity, the models can capture the main features of particle coating for ALD, at least in the case of batch fluidized bed reactors.